If a 387-mL sample of neon gas is heated from 21°C to 71°C at constant pressure, what will be the volume of the sample at the higher temperature?
Answer #1

The given information:

The volume of Gas = 387 ml

Initial temperature = 21 + 273 = 294 k

Final temerature = 71 + 273 = 344 k

The volume of the sample at a higher temperature = 452.81 ml.